# Methodology for Fine-Grain GPU Power Visibility and Insights

### v1

@@ V-B FinGraV Methodology Evaluation
Before we present AI power profiles, we first begin with an evaluation of the key tenets of FinGraV methodology namely: (a) the benefit of CPU-GPU time sync, (b) the effect of kernel execution time binning, and (c) resiliency to #runs executed. To do so, we focus on multiple power profiles for a compute-bound (CB) 4K GEMM kernel (hence referred to as CB-4K-GEMM for simplicity) depicted in Figure 5 with and without these techniques. We depict in the figure multiple runs of a CB-4K-GEMM and multiple executions within a run (about 25). We have time for a run on the x-axis and total power profiled on the y-axis. [^1] Note that, in this work, we will focus on relative power data only (not absolute power numbers). Additionally, using vertical line markers, we separate the warmup, SSE (steady-state execution) and SSP (steady-state power) executions on the graph (Section IV-A).
Benefits of CPU-GPU Time Sync: We discussed in Section IV-A the importance of syncing CPU-GPU time while using a GPU-based power logger for events triggered by the CPU. Figure 5 shows the benefit of this time synchronization for CB-4K-GEMM by comparing the unsynchronized (red) power profile to the synchronized (blue) power profile. As shown, the synchronized profile captures the gradual rise in power as GPU moves from idle state to executing the kernel (warm-up executions, to SSE, then finally to SSP) while the unsynchronized profile misses this ramp and fails to align power changes with appropriate executions in a run. [^2] Figure 5, and subsequent profiles, depict a single SSP kernel execution, which is the last execution of a given kernel in a given run (Section IV-A).
@@
TABLE II: FinGraV Profiling Insights and Recommendations for Future Hardware/Software.
# Takeaway Recommendation Section
1 Similar kernel execution times can manifest very different power profiles. Considerable smarts in power management (firmware) are necessary to be power proportional on GPUs. Section V-C1
2 Total power scales with work done and different GPU components get stressed based on computation behavior. Concurrency benefits can be reaped via stitching together kernels with complementary power profiles (given power headroom). Section V-C2
3 Compute-light kernels and compute-heavy kernels show similar compute power. Techniques to lower compute power for compute-light kernels are necessary. Section V-C2
4 Power for power-light kernels (memory-bound GEMVs and compute-light GEMMs) gets affected by kernels preceding them while power-heavy kernels are not affected. Power management mechanisms which lower the power penalty incurred by power-light kernels when they follow power-heavy kernels are necessary. Section V-C3
5 When possible, at the software-level (scheduler) or at the algorithm design, prioritizing low-power kernels in scheduling can help. Section V-C3
@@ V-C GEMM Profiles and Insights
We discuss our GEMM power profiles (compute-bound/CB and memory-bound/MB) and learnings in this section. We start with general power trends we observe, follow that with component-level (Section II-A) comparative analysis, and finally discuss power behavior in the presence of interleaved GEMM executions. Along the way, we make several key observations and recommendations for future hardware/software to optimize GPU power (all summarized in Table II).
@@ V-C1 Power Trends
We first begin with the general question of power trends manifested. To do so, we depict the power profile (total and XCD component) for 8K compute-bound GEMM (CB-8K-GEMM) over 200 runs with eleven executions per run in Figure 6. We make several observations here. First, the power rises for initial executions and then it drops till it reaches SSE profile. Finally, the power slightly increases till it reaches SSP profile at the eleventh execution after which it shows little variation. Also, we note that the total power is dominated by XCD power for such CB GEMMs.
We surmise that for a compute-heavy GEMM kernel such as CB-8K-GEMM, the first few executions considerably stress power, invoking the power management firmware to drop the power. Such behavior has been documented before [18]. We also observe that the best execution time, which happens for SSE execution and beyond, shows lower power than these initial executions. Finally, the SSP execution shows slightly higher power than the SSE execution.
We compare these observations to the power profile for CB-2K-GEMM (Figure 7) which is considerably less compute-heavy in comparison to CB-8K-GEMM (based on op-to-byte ratio). We observe that power starts low for initial executions before rising considerably for the SSP execution (40th execution in case of CB-2K-GEMM). Recall that all executions between SSE and SSP have similar execution times. Overall, the two key power trends we observe are (a) a sharp rise followed by a drop to a steady-state and (b) a gradual rise to a steady-state.
The above observations lead us to our first key takeaway (highlighted in Table II) that similar kernel execution times (SSE through SSP) can manifest very different power profiles. This in turn leads us to make our first recommendation that considerable smarts in power management (firmware) are necessary to ensure that the lowest power profile is exercised for a given computation.
@@ V-C2 Component Comparative Analysis
Next, we compare component-level power across all three CB GEMMs and three MB GEMVs. The relative powers for total power and the component-level powers for XCD, IOD, and HBM is depicted in Figure 8. Note that we use SSP power profiles of these kernels to plot these graphs. That said, SSE profiles show similar trends with one exception which we highlight below. Finally, for better visibility, we also show linear regression lines for all power profiles.
@@
We make several observations here. All CB GEMMs show considerably higher total and XCD power versus MB GEMVs. This makes sense as CB GEMMs incur heavy compute and also data movement, while MB GEMVs are data movement heavy only. [^3] As we repeatedly execute kernels, data movement is heavily biased toward on-chip data movement for our executions. Amongst CB GEMMs, CB-8K-GEMM has slightly higher total/XCD power. On the GEMV front, we see a drop in total power going from 8K-GEMV to 2K-GEMV. Unlike total/XCD power, where CB GEMMs dominate, MB-8K-GEMV does stress IOD power. Finally, as discussed above, as our data movement is biased towards on-chip data movement (repeated executions), both CB/MB GEMM kernels stress HBM power similarly (except CB-8K-GEMM, which has the highest HBM power, as its considerably large input sizes stress the on-chip caches the most).
Figure 9: Total power comparison for different interleaved GEMM and GEMV experiments.
@@
The above observations lead us to our second key takeaway (Table II) that total power generally scales with work done (e.g., CB > MB) with different components getting stressed based on computation behavior (e.g., CB stress XCD power, MB can stress IOD power, etc.). This takeaway in turn leads to our second recommendation that, should it be possible, we should stitch together kernels with complementary power profiles at an algorithmic/software level. This, with enough power headroom, will allow us to reap the benefits of concurrent execution. An example of this for AI workloads is concurrent execution of MB attention kernel with CB fully-connected layers [19].
@@
Another takeaway we point to is that all CB GEMMs are in the ballpark of each other when it comes to XCD power. That said, compute throughput calculated using algorithmic ops and execution times shows that CB-2K-GEMM has about half the compute utilization in comparison to CB-4K/8K-GEMM. This leads us to our third recommendation that techniques which lower the compute power for compute-light kernels (e.g., CB-2K-GEMM with comparatively lower compute utilization) are necessary. Note that, unlike XCD power, we observe that the IOD power tracks well with LLC bandwidth, underscoring the importance of focusing on the XCD component.
Finally, all of the above component-level observations also hold with SSE profiles, albeit at lower magnitude and with one exception. A compute-light kernel, such as CB-2K-GEMM, manifests power similar to MB GEMMs when it comes to its SSE profile.
@@ V-C3 Interleaved Kernels Analysis
We see something similar for MB kernels in the graph on the right in Figure 9. MB kernels, being power-light, are affected by kernels that precede them. Comparing the SSP profile for MB-8K-GEMV, when other MB kernels (MB-4K/2K-GEMV) are interleaved before it, shows that MB-8K-GEMV consumes less power in comparison to its SSP profile (MB–>8K_gemv). On the other hand, MB-4K-GEMV shows more than its SSP power when CB kernels (CB-8K/4K-GEMM) are run before it (CB–>4K_gemv).
The above observations lead us to our fourth key takeaway that power-light kernels (MB GEMVs and compute-light GEMMs) are affected by kernels preceding them when it comes to their power consumption. This takeaway leads to two recommendations from us. First, power management mechanisms that lower the power penalty incurred by power-light kernels when they follow power-heavy kernels are necessary. This can include (but not be limited to) pro-active fine-grain voltage/frequency modulation mechanisms in power management firmware. That said, further thought is necessary in this regard on the appropriate management of power swings (high-to-low, low-to-high) as these can have non-trivial secondary effects [20]. Another recommendation is that, when possible and when we are not limited by data dependency, prioritizing the scheduling of power-light kernels over power-heavy kernels can be practiced.
@@ V-D Communication Profiles and Insights
We next focus on power analysis for communication kernels. Recall that we profile two widely used AI communication kernels: all-gather (AG) and all-reduce (AR). Further, we exercise both latency-bound/LB (64KB/128KB) and bandwidth-bound/BB scenarios (512MB/1GB) for both kernels.
At a high level, unlike GEMM kernels, we don’t see SSE/SSP differentiation for communication kernels. As such, we focus on SSE profiles for communication kernels. Further, Figure 10 presents relative powers (total, XCD, IOD, and HBM) for all eight communication kernels. We also plot CB-8K-GEMM in the same graph for comparing GEMM and communication kernel power profiles.
We make several observations here. With regards to XCD power, CB-8K-GEMM has much higher power than communication kernels which is expected. That said, when it comes to the total power, BB communication kernels fall somewhere in the middle of LB communication and CB-GEMM kernels. This can be explained by the considerably higher IOD and HBM power incurred by BB communication kernels.
Based on the comparison of communication and computation-heavy (GEMM) kernels, our takeaway is similar to takeaway #2 in Table II that the heterogeneous power profiles manifested can be exploited for efficient concurrent execution should there be enough power headroom.
@@ VI Discussion
FinGraV with External Power loggers: In our work, we harness an internal 1ms power logger (Section IV-A). We believe that the general FinGraV methodology we have can be applied to external power loggers available on AMD platforms such as amd-smi [14]. The key tenets of FinGraV (careful CPU-GPU time synchronization, kernel execution time binning, and power profile differentiation) are all equally relevant even with these external power loggers. All of this said, the resultant power profiles will heavily depend on the power information these external loggers report. As an example, since these loggers may report average power, any averaging done can impact the power profiles that FinGraV methodology produces. We leave further augmenting FinGraV for these tools to future work.
We believe that, in addition to FinGraV methodology, a key contribution of our work is indeed the FinGraV profiles for key AI operations of GEMMs and communication kernels. Additionally, another key contribution of our work is the recommendations we have for future software/hardware to optimize power. All of these are immensely valuable in designing power-optimized future accelerators.
Outlier Executions: We discussed in Section IV-A that we focus FinGraV profiles on the common kernel execution time and discard any outliers. While we believe that prioritizing power analysis for the common case is the right first step, understanding the power behavior of the outlier executions is also important. One way to attain FinGraV profiles for outlier executions is to employ FinGraV methodology and focus on collecting profiles for a specific outlier execution time and discarding the rest (that is changing step-5 in Section IV-B). While doable, this can be costly as more #runs can be necessary to create a large enough bin belonging to this specific outlier execution time. Another strategy could be to break down kernel execution into phases when possible and assess if this lowers variation in each phase (as compared to variation at the kernel-level). As an example, with GPU kernels, wherein each kernel launches multiple workgroups, the kernel can be artificially terminated after half the number of workgroups are completed and each half of the execution can be studied separately. We leave such investigations to future work.
@@ VII Related Work
Power/Energy Measurement Methodologies: Prior work measured GPU power using vendor tools and interfaces [14, 21] to access native on-board telemetry readings. Tools, such as Variorum [22], harness these interfaces to provide a vendor-neutral library for power measurement across different hardware from multiple vendors. Such efforts to standardize the power measurement interface, along with exposing more telemetry and control, are critical for exascale systems as highlighted by national labs and hyperscalers [23, 24]. To sample power at a higher rate compared to vendor tools, other work physically measure power using external power meters [25]. Using these meters, prior work assessed the existing on-board power measurement and found discrepancies in the reported power, highlighting the importance of using a high-fidelity methodology for measuring and reporting power [26, 27, 28, 29, 30]. Finally, researchers built simulators and statistical models to provide fine-grain power and energy estimates [31, 32, 33, 34, 35, 36].
In this work, we focus on native on-board power measurements, instead of modeling, as the GPUs are getting more complex, pushing both compute and memory. With its fine-grain power profiles, FinGraV can be used to improve the fidelity of the power models. As discussed in Section VI, FinGraV can work with existing public power loggers (with the caveats discussed) and can complement prior power measurement methods by providing a step-by-step methodology to unlock fine-grain visibility for the evaluated kernels. This way FinGraV can be agnostic to the varying sampling rate of the power measurement interfaces and meters. As the kernel execution gets shorter, FinGraV key principles will be more critical to collect sane fine-grain power profiles.
@@
We harnessed FinGraV to collect fine-grain power profiles for sub-ms GEMM/V kernels and communication collectives, key AI primitives, running on AMD MI300X. These power profiles, across the sub-components of MI300X, unveiled insights related to the different power behaviors of the evaluated kernels during standalone runs and when interspersed between other kernels. This led to the key power-aware recommendations to optimize both future GPU software and hardware. Overall, using FinGraV fine-grain profiles, researchers can uncover challenges and energy bloats to better design GPU power managers to improve the energy efficiency of standalone GPUs and across large-scale clusters.

### v2

@@ V-B FinGraV Methodology Evaluation
Before we present AI power profiles, we first begin with an evaluation of the key tenets of FinGraV methodology namely: (a) the benefit of CPU-GPU time sync, (b) power profile differentiation benefit, (c) the effect of kernel execution time binning, and (d) resiliency to #runs executed. To do so, we focus on multiple power profiles for a compute-bound (CB) 4K GEMM kernel (hence referred to as CB-4K-GEMM for simplicity) depicted in Figure 5 with and without these techniques. In the figure, we depict multiple runs of a CB-4K-GEMM and multiple executions within a run. We have time for a run on the x-axis and total power profiled on the y-axis. [^1] In this work, we focus on relative power data (not absolute numbers). [^2] The total power, on the y-axis, is the voltage regulator output power. Additionally, using vertical line markers, we separate the warmup, SSE (steady-state execution) and SSP (steady-state power) executions on the graph (Section IV-A).
@@
Benefits of CPU-GPU Time Sync: We discussed in Section IV-A the importance of syncing CPU-GPU time while using a GPU-based power logger for events triggered by the CPU. Figure 5 shows the benefit of this time synchronization for CB-4K-GEMM by comparing the unsynchronized (red) power profile to the synchronized (blue) power profile. As shown, the synchronized profile captures the gradual rise in power as GPU moves from idle state to executing the kernel (warm-up executions, to SSE, then finally to SSP) while the unsynchronized profile misses this ramp and fails to align power changes with appropriate executions in a run.
Accurate Power Profiles with Power Profile Differentiation: Figure 5 shows that SSE and SSP profiles differ considerably. Consequently, assuming SSE profile as the kernel’s power profile can lead to up to 36% error in power (and hence energy) measurement.
@@ V-C GEMM Profiles and Insights
We discuss our GEMM power profiles (compute-bound/CB and memory-bound/MB) and learnings in this section. We start with general power trends we observe, follow that with component-level (Section II-A) comparative analysis, and finally discuss power behavior in the presence of interleaved GEMM executions. Along the way, we make several key observations, provide guidance for accurate power measurements and make recommendations for future hardware/software to optimize GPU power (all summarized in Table II).
@@ V-C1 Power Trends
We first begin with the general question of power trends manifested. To do so, we depict the power profile (total and XCD component) for 8K compute-bound GEMM (CB-8K-GEMM) over 200 runs with multiple executions per run in Figure 6. We make several observations here. First, the power rises for initial executions and then it drops till it reaches SSE profile. Finally, the power slightly increases till it reaches SSP profile at the last execution after which it shows little variation (Section IV-A).
We observe that for a compute-heavy GEMM kernel such as CB-8K-GEMM (high op-to-byte ratio), the first few executions considerably stress power, invoking the power management firmware to throttle frequency [5] in order to manage power excursions. We also observe that the best execution time, which happens for SSE execution and beyond, shows lower power than these initial executions. Finally, the SSP execution shows slightly higher power than the SSE execution.
We compare these observations to the power profile for CB-2K-GEMM (Figure 8) which is considerably less compute-heavy in comparison to CB-8K-GEMM (based on op-to-byte ratio). We observe that power starts low for initial executions before rising considerably for the SSP execution. Recall that all executions between SSE and SSP have similar execution times.
@@
Based on above two power profiles, overall, the two key power trends we observe are (a) a sharp rise followed by a drop to a steady-state and (b) a gradual rise to a steady-state. This provides credence to FinGraV power profile differentiation strategy that eventually power stabilizes giving a time-series view of average power at different points in kernel execution time (SSP profile). Further, as execution time of CB-8K-GEMM is longer than power averaging window (1ms) while CB-2K-GEMM is much shorter, we observe that the resultant spread between SSE and SSP power, and hence power/energy measurement error, is much higher for CB-2K-GEMM (80% vs. 20% for CB-8K-GEMM).
The above observations lead us to our first key takeaway (highlighted in Table II) that similar kernel execution times (SSE through SSP) can manifest very different power profiles. Further, not factoring this variation can lead to a measurement error of as high as 80% depending on relative magnitudes of kernel execution time and power averaging time window of underlying power logger. This also leads to our first power measurement guidance that power profile differentiation (SSE, SSP) as employed by FinGraV is very crucial for accurate power measurements.
@@ V-C2 Component Comparative Analysis
Next, we compare component-level power across all three CB GEMMs and three MB GEMVs. Figure 7 depicts, in a relative manner, the total and component-level powers (XCD, IOD, HBM). Note that we use SSP power profiles of these kernels to plot these graphs for as discussed above SSP profile represents the true time-series view of average power for different points in kernel execution. Finally, for better visibility, we also show linear regression lines for all power profiles.
@@
We make several observations here. All CB GEMMs show considerably higher total and XCD power versus MB GEMVs. This makes sense as CB GEMMs incur heavy compute and also data movement, while MB GEMVs are data movement heavy only. [^3] As we repeatedly execute kernels, data movement is heavily biased toward on-chip data movement for our executions. Amongst CB GEMMs, CB-8K-GEMM has slightly higher total/XCD power. On the GEMV front, we see a drop in total power going from 8K-GEMV to 2K-GEMV. Unlike total/XCD power, where CB GEMMs dominate, MB-8K-GEMV does stress IOD power. Finally, as discussed above, as our data movement is biased towards on-chip data movement (repeated executions), both CB/MB GEMM kernels stress HBM power similarly. The exception is CB-8K-GEMM, which has the highest HBM power, as its considerably large input sizes stress the on-chip caches the most.
TABLE II: FinGraV Profiling Insights, Power Measurement Guidance and, Recommendations for Future Hardware/Software.
# Takeaway Power Measurement Guidance/Recommendation Section
1 Similar kernel execution times can manifest very different power profiles depending on relative magnitudes of kernel execution time and power averaging time-window of underlying power logger. Measurement guidance-1: FinGraV power profile differentiation is crucial to deduce a kernel’s power profile without which high measurement errors (as high as 80%) can manifest. Section V-C1
2 Total power scales with work done and different GPU components get stressed based on algorithmic nature of underlying computation. Recommendation-1: Available power headroom can be fully utilized by concurrently executing computations with complementary algorithmic and hence complementary power profiles. Section V-C2
3 Compute-heavy kernels are dominated by XCD component power. Recommendation-2: Techniques that optimize XCD component power should be prioritized to optimize overall total power for compute-heavy kernels. Section V-C2
4 Compute-light kernels and compute-heavy kernels show similar XCD component power. Recommendation-3: Techniques to attain GPU power proportionality are necessary for compute-light kernels especially for XCD component. Section V-C2
5 Power for certain kernels (memory-bound GEMVs and compute-light GEMMs) gets affected by kernels preceding them while for other kernels (compute-heavy GEMMs) it is not affected. Measurement guidance-2: Isolated executions are necessary to assess kernel’s power draw when its execution time is shorter than power averaging time-window of underlying power logger. Section V-C3
@@
The above observations lead us to our second key takeaway (Table II) that total power generally scales with work done (e.g., CB > MB) with different components getting stressed based on the underlying algorithm for a given computation (e.g., CB stress XCD power, MB can stress IOD power, etc.). That is, algorithmic behavior (op-to-byte) is, to a first order, a good indicator of component-level GPU power consumption. This takeaway in turn leads to our first recommendation that, should it be possible, we should stitch together kernels with complementary power profiles at an algorithmic/software level. This, with enough power headroom, will allow us to reap the benefits of concurrent execution. An example of this for AI workloads is concurrent execution of MB attention kernel with CB fully-connected layers [18].
Further note that, combining the data in Figure 6 and Figure 8, it is clear that total power and XCD power are very close to each other for CB GEMMs. This, together with low IOD/HBM power depicted in Figure 7, leads to our third key takeaway that XCD component is the most dominant component for CB GEMMs which leads, by Amdahl’s law, to our second recommendation that techniques that focus on optimizing XCD power are crucial to optimize overall power for CB computations.
Another takeaway we point to is that all CB GEMMs are in the ballpark of each other when it comes to XCD power. That said, compute throughput calculated using algorithmic ops and execution times shows that CB-2K-GEMM has about half the compute utilization in comparison to CB-4K/8K-GEMM. Note that, unlike XCD power, we observe that the IOD power tracks well with LLC bandwidth. This leads us to our third recommendation that GPU power proportionality, that is, expending power commensurate to rate of work, needs further investigation especially with a focus on XCD component for compute-light kernels (e.g., CB-2K-GEMM with comparatively lower compute utilization). Of particular focus can be different software execution schedules which are performance-iso but lead to different (hopefully lower) power profiles.
Figure 9: Total power comparison for different interleaved GEMM and GEMV.
@@ V-C3 Interleaved Kernels Analysis
We see something similar for MB kernels in the graph on the right in Figure 9. MB kernels are also affected by kernels that precede them. Comparing the SSP profile for MB-8K-GEMV, when other MB kernels (MB-4K/2K-GEMV) are interleaved before it, shows that MB-8K-GEMV consumes less power in comparison to its SSP profile (MB–>8K_gemv). On the other hand, MB-4K-GEMV shows more than its SSP power when CB kernels (CB-8K/4K-GEMM) are run before it (CB–>4K_gemv).
The above observations lead us to our fifth key takeaway that certain kernels (MB GEMVs and compute-light GEMMs) are affected by kernels preceding them when it comes to their power consumption. The primary reason for this behavior is that the execution time of these kernels is considerably shorter than power averaging time window of the power logger we use. As such, their measured power profiles are an average of power manifested by preceding kernels and the kernel of interest. This in turn leads to our second measurement guidance that for kernels whose execution time is lower than power averaging time-window, isolated executions are necessary to assess the true kernel power draw. Note that, with an instantaneous power sampler, FinGraV methodology can assess power for kernels regardless of their execution time and regardless of run setup (isolated or interleaved executions).
@@ V-D Communication Profiles and Insights
We next focus on power analysis for communication kernels. Recall that we profile two widely used AI communication kernels: all-gather (AG) and all-reduce (AR). Further, we exercise both latency-bound/LB (64KB/128KB) and bandwidth-bound/BB scenarios (512MB/1GB) for both kernels. Figure 10 presents relative powers (total, XCD, IOD, and HBM) for all eight communication kernels. We also plot CB-8K-GEMM in the same graph for comparing GEMM and communication kernel power profiles.
We make several observations here. With regards to XCD power, CB-8K-GEMM has much higher power than communication kernels, which is expected. That said, when it comes to the total power, BB communication kernels fall somewhere in the middle of LB communication and CB-GEMM kernels. This can be explained by the considerably higher IOD and HBM power incurred by BB communication kernels.
Based on the comparison of communication and computation-heavy (GEMM) kernels, our takeaway is similar to takeaway #2 in Table II and consequently our resultant recommendation is also the same in that the heterogeneous power profiles manifested can be exploited for efficient concurrent execution should there be enough power headroom (e.g., latency-bound communication in parallel with any other computation, etc.).
@@ VI Discussion
FinGraV with External Power loggers: In our work, we harness a power logger available internally at AMD on MI300X that provides average power samples for multiple instantaneous power readings in last 1ms (Section IV-A). We believe that the general FinGraV methodology we have can be applied to external power loggers available on AMD platforms such as amd-smi [14]. The key tenets of FinGraV (careful CPU-GPU time synchronization, kernel execution time binning, and power profile differentiation) are all equally relevant even with these external power loggers. All of this said, the resultant power profiles will heavily depend on the power information these external loggers report. As an example, since these loggers may report average power over longer time-windows, any such averaging done can impact the power profiles that FinGraV methodology produces. We leave further augmenting FinGraV for these tools to future work.
We believe that, in addition to FinGraV methodology, a key contribution of our work is indeed the power profiles for key AI operations of GEMMs and communication kernels that we make available. These profiles provide both an accurate and deeper (different components such as XCD, IOD, HBM) view of power consumption on a state-of-the-art MI300X GPU which led to some targeted recommendations for power optimizations (Table II). Finally, we also provided power measurement guidance (Table II), which will aid researchers in avoiding common pitfalls and steep errors that can happen with power (and hence energy) measurements on high-end GPUs (as high as 80% error). All of these are immensely valuable in designing power-optimized future accelerators.
Outlier Executions: We discussed in Section IV-A that we focus FinGraV profiles on the common kernel execution time and discard any outliers. While we believe that prioritizing power analysis for the common case is the right first step, understanding the power behavior of the outlier executions is also important. One way to attain FinGraV profiles for outlier executions is to employ FinGraV methodology and focus on collecting profiles for a specific outlier execution time and discarding the rest (that is changing step-6 in Section IV-B). While doable, this can be costly as more #runs can be necessary to create a large enough bin belonging to this specific outlier execution time. Another strategy could be to break down kernel execution into phases when possible and assess if this lowers variation in each phase (as compared to variation at the kernel-level). As an example, with GPU kernels, wherein each kernel launches multiple workgroups, the kernel can be artificially terminated after half the number of workgroups are completed and each half of the execution can be studied separately. We leave such investigations to future work.
@@ VII Related Work
Power/Energy Measurement Methodologies: Prior work measured GPU power using vendor tools and interfaces [14, 19] to access native on-board telemetry readings. Tools, such as Variorum [20], harness these interfaces to provide a vendor-neutral library for power measurement across different hardware from multiple vendors. Such efforts to standardize the power measurement interface, along with exposing more telemetry and control, are critical for exascale systems as highlighted by national labs and hyperscalers [21, 4]. To sample power at a higher rate compared to vendor tools, other work physically measure power using external power meters [22]. Using these meters, prior work assessed the existing on-board power measurement and found discrepancies in the reported power, highlighting the importance of using a high-fidelity methodology for measuring and reporting power [23, 24, 25, 26, 6]. Concurrent to our work, Yang et al. [6] did a comprehensive analysis of non-AMD GPUs and identified similar power measurement guidance as we do (e.g., skip/shift mechanism they employ is equivalent to SSE/SSP power profile differentiation we employ). However, they do not evaluate AMD GPUs and further, they do not focus on power optimization recommendations. Finally, researchers built simulators and statistical models to provide fine-grain power and energy estimates [27, 28, 29, 30, 31, 32].
In this work, we focus on native on-board power measurements, instead of modeling, as the GPUs are getting more complex, pushing both compute and memory throughputs and hence power. With its fine-grain power profiles, FinGraV can be used to improve the fidelity of the power models. As discussed in Section VI, FinGraV can work with existing public power loggers (with the caveats discussed) and can complement prior power measurement methods by providing a step-by-step methodology to unlock fine-grain visibility for the evaluated kernels. This way FinGraV can be agnostic to the varying sampling rate of the power measurement interfaces and meters. As the kernel executions get shorter, FinGraV key principles will be more critical to collect sane fine-grain power profiles.
@@
We harnessed FinGraV to collect fine-grain power profiles for sub-ms GEMM/V kernels and communication collectives, key AI primitives, running on AMD MI300X. These power profiles, across the sub-components of MI300X, unveiled insights related to the different power behaviors of the evaluated kernels. Using FinGraV fine-grain profiles, researchers can uncover opportunities to design better GPU power managers to improve the energy efficiency of both, standalone GPUs and large-scale GPU clusters.